OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is odd and one die shows a 5? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

table 6x6 will help u

OpenStudy (anonymous):

i dont know how to use one.

OpenStudy (anonymous):

1 2 3 4 5 6 1 2 3 4 5 6 7 2 3 4 5 6 7 8 3 4 5 6 7 8 9 4 5 6 7 8 9 10 5 6 7 8 9 10 11 6 7 8 9 10 11 12

OpenStudy (anonymous):

sum of the dice is odd = 18/36 = 1/2

OpenStudy (anonymous):

it was 1/6 ..

OpenStudy (anonymous):

hmm: 3 = 2 5 = 4 7 = 6 9 = 4 11 = 2 ---- 18

OpenStudy (anonymous):

idk my online quiz said it was 1/6..thnx though